# Certain finite groups whose commuting conjugacy class graph satisfy Hansen-Vukičević conjecture

### v1

@@ 2.2. Zagreb indices of CCC-graphs of groups with a given central quotient
We have $\mathcal{CCC}(G)=K_{\frac{x(q-1)}{p}}\sqcup K_{x(p-1)}$, where $x=|Z(G)|$. As such, $|v(\mathcal{CCC}(G))|=\frac{x(q-1)}{p}+x(p-1)=\frac{x}{p}(p^{2}+q-p-1)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{x(q-1)}{p}}{2}+\binom{x(p-1)}{2}=\frac{x}{2p^{2}}(p^{2}x+q^{2}x+p^{4}x+x-2qx-pq-2p^{3}x-p^{3}+p^{2}+p)$. Also, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\frac{x(q-1)}{p}\left(\frac{x(q-1)}{p}-1\right)^{2}+x(p-1)\left(x(p-1)-1\right)^{2}$
$=\frac{x}{p^{3}}\left(q^{3}x^{2}-3q^{2}x^{2}+3qx^{2}+p^{6}x^{2}-3p^{5}x^{2}+3p^{4}x^{2}-p^{3}x^{2}-x^{2}-2pq^{2}x\right.$
$M_{2}(\mathcal{CCC}(G))$ $=\frac{x(q-1)}{p}\frac{\left(\frac{x(q-1)}{p}-1\right)^{3}}{2}+x(p-1)\frac{\left(x(p-1)-1\right)^{3}}{2}$
$=\frac{x}{2p^{4}}\left(q^{4}x^{3}-4q^{3}x^{3}+6q^{2}x^{3}-4px^{3}-p^{3}q-4p^{7}x^{3}+6p^{6}x^{3}-4p^{5}x^{3}+3px^{2}\right.$
$\,\,\,\left.+p^{4}x^{3}-6p^{2}qx+3p^{6}x-6p^{5}x+3p^{4}x+3p^{2}x+x^{3}+p^{8}x^{3}-p^{5}+p^{4}+p^{3}\right).$
Therefore
$\qquad\qquad=\frac{f(p,q,x)}{p^{2}(p^{2}+q-p-1)(p^{2}x+q^{2}x+p^{4}x+x-2qx-pq-2p^{3}x-p^{3}+p^{2}+p)}$
$\qquad\qquad:=\frac{f(p,q,x)}{g(p,q,x)},$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+10p^{5}x^{2}-3p^{4}x^{3}-10p^{4}x^{2}+6p^{3}x^{2}-2p^{2}x^{2}.$
On factorization, we have $f(p,q,x)=px^{2}(p-1)(q-1)(p+q-p^{2}-1)^{2}(p^{2}x+qx-x-2p-px)$. Suppose that $p^{2}x+qx-x-2p-px=p(x(p-1)-2)+x(q-1):=S_{1}+S_{2}$, where $S_{1}=p(x(p-1)-2)$ and $S_{2}=x(q-1)>0$ for all prime $q$ and $x\geq 1$. Clearly, $S_{1}\geq 0$ for all prime $p\geq 3$ and $x\geq 1$. Now, $f(2,q,1)=2(q-1)(q-3)^{2}(q-3)\geq 0$ for all $q\geq 3$. Also, $g(p,q,x)=\frac{2p^{5}}{x^{2}}\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. Therefore, $\frac{f(p,q,x)}{g(p,q,x)}\geq 0$ and hence the result follows. ∎
@@
In all the cases, we have $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}>\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ and hence the result follows.
@@
Also, $|v(\mathcal{CCC}(G))|=\frac{x(p^{2}-1)}{p}+p^{2}\times\frac{x(p-1)}{p^{2}}=x(p-1)(p+2)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{x(p^{2}-1)}{p}}{2}+p^{2}\binom{\frac{x(p-1)}{p^{2}}}{2}=\frac{x(p-1)}{2p^{2}}(p^{3}x+p^{2}x-2p^{2}-2x-p)$. Therefore
@@
$\,=\frac{1}{g(p,x)}\left(p^{10}x^{3}+2p^{9}x^{3}-3p^{9}x^{2}-3p^{8}x^{3}-5p^{8}x^{2}-8p^{7}x^{3}+3p^{7}x+3p^{6}x^{3}+17p^{6}x^{2}\right.$
$\quad\quad\quad\quad\quad\left.+11p^{5}x^{3}-9p^{5}x+3p^{8}x-15p^{4}x^{2}-8p^{3}x^{3}-p^{6}-2p^{7}+3px^{3}+3p^{3}x^{2}\right.$
$\quad\quad\quad\quad\quad\left.-2p^{2}x^{3}+p^{4}x^{3}+3p^{2}x^{2}+7p^{7}x^{2}-8p^{5}x^{2}-3p^{6}x+6p^{3}x+2p^{5}+p^{4}+px^{2}\right)$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.+p^{6}x^{2}(x-3)+3p^{5}x+p^{2}x(5x^{2}-4x-6)+p^{4}(4x^{2}+3x-2)\right.\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.+p^{3}(x^{3}+12x^{2}-1)\right)\right)$
@@
Also, $|v(\mathcal{CCC}(G))|=(p^{2}+p+1)\frac{x(p-1)}{p^{2}}=\frac{x(p^{3}-1)}{p^{2}}$ and $|e(\mathcal{CCC}(G))|=(p^{2}+p+1)\binom{\frac{x(p-1)}{p^{2}}}{2}$ $=\frac{x(p^{3}-1)}{2p^{4}}(px-x-p^{2})$. Therefore
$\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{p^{2}x^{2}-2px^{2}+x^{2}-2p^{3}x+2p^{2}x+p^{4}}{p^{4}}$
and
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}$ $=\frac{p^{3}x^{3}-3p^{2}x^{3}+3px^{3}-x^{3}-3p^{4}x^{2}+6p^{3}x^{2}-3p^{2}x^{2}+3p^{5}x-3p^{4}x-p^{6}}{p^{4}(px-x-p^{2})}$
$=\frac{(px-x-p^{2})(p^{2}x^{2}-2px^{2}+x^{2}-2p^{3}x+2p^{2}x+p^{4})}{p^{4}(px-x-p^{2})}$
$=\frac{p^{2}x^{2}-2px^{2}+x^{2}-2p^{3}x+2p^{2}x+p^{4}}{p^{4}}=\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}.$
@@
Also, $|v(\mathcal{CCC}(G))|=\frac{x(p^{2}-1)}{p}+kp\frac{x(p-1)}{p^{2}}+(p-k)\frac{x(p-1)}{p}=\frac{x(p-1)}{p^{2}}(2p^{2}+p)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{x(p^{2}-1)}{p}}{2}+kp\binom{\frac{x(p-1)}{p^{2}}}{2}+(p-k)\binom{\frac{x(p-1)}{p^{2}}}{2}=\frac{x(p-1)}{2p^{4}}(p^{5}x+2p^{4}x-2p^{3}x-2p^{4}-p^{2}x-p^{3}+2p^{2}kx-kpx-kp^{3}x)$. Therefore
$\qquad=\frac{f_{1}(p,k,x)}{p^{4}(2p^{2}+p)(p^{5}x+2p^{4}x-2p^{3}x-2p^{4}-p^{2}x-p^{3}+2p^{2}kx-kpx-kp^{3}x)}$
$\qquad:=\frac{f_{1}(p,k,x)}{g_{1}(p,k,x)},$
@@
Also, $|v(\mathcal{CCC}(G))|=(kp+1)\frac{x(p-1)}{p^{2}}+(p+1-k)\frac{x(p-1)}{p}=\frac{x(p-1)}{p^{2}}(p^{2}+p+1)$ and $|e(\mathcal{CCC}(G))|=(kp+1)\binom{\frac{x(p-1)}{p^{2}}}{2}+(p+1-k)\binom{\frac{x(p-1)}{p}}{2}=\frac{x(p-1)}{2p^{4}}(2kp^{2}x-kpx+px-x-p^{2}+p^{4}x+p^{4}-p^{2}x-p^{3}-kp^{3}x)$. Therefore
$\!\!\!\!\!=\frac{f_{2}(p,k,x)}{p^{4}(p^{2}+p+1)(2kp^{2}x-kpx+px-x-p^{2}+p^{4}x+p^{4}-p^{2}x-p^{3}-kp^{3}x)}:=\frac{f_{2}(p,k,x)}{g_{2}(p,k,x)},$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+2kp^{3}x^{2}-2p^{3}x^{2}.$
On factorization, we have $f_{2}(p,k,x)=px^{2}(p-1)^{4}(p+1-k)(kp+1)(p^{2}x-x-2p^{2})$. For all $p\geq 2,1\leq k\leq p$ and $x\geq p^{2}$, we have $p+1-k>0$ and $x(p^{2}-1)>2p^{2}$. As such, $f_{2}(p,k,x)>0$. Also, $g_{2}(p,k,x)=\frac{2p^{10}}{x^{2}(p-1)^{2}}\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. Therefore, $\frac{f_{2}(p,k,x)}{g_{2}(p,k,x)}>0$.
@@ 3. Concluding remark
As a consequence it was also seen that if $G$ is any non-abelian group of order $p^{3}$ or $p^{4}$, where $p$ is any prime, then $\mathcal{CCC}(G)$ satisfies Hansen-Vukičević conjecture. Using Theorem 2.8, Theorem 2.9 and the characterizations of $4$-centralizer (“$G$ is a $4$-centralizer finite group if and only if $\frac{G}{Z(G)}\cong\mathbb{Z}_{2}\times\mathbb{Z}_{2}$[3, Theorem 2]), $5$-centralizer (“$G$ is a $5$-centralizer finite group if and only if $\frac{G}{Z(G)}\cong\mathbb{Z}_{3}\times\mathbb{Z}_{3}$ or $D_{2\times 3}$[3, Theorem 4]) finite groups we have the following result.
@@
Further, if $\Pr(G)=\frac{5}{14},\frac{2}{5},\frac{11}{27},\frac{1}{2}$ or $\frac{5}{8}$ then $\frac{G}{Z(G)}\cong D_{2\times 7},D_{2\times 5},D_{2\times 4},D_{2\times 3}$ or $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ (see [24, pp. 246] and [21, pp. 451]). Therefore, the following result follows in view of Theorem 2.8 and Theorem 2.9.
@@
In our study we have not found any group $G$ such that $\mathcal{CCC}(G)$ does not satisfy Hansen-Vukičević conjecture. We conclude this paper with the following question.
@@
Acknowledgement. The first author is thankful to Council of Scientific and Industrial Research for the fellowship (File No. 09/0796(16521)/2023-EMR-I).
Declarations
  • Funding: No funding was received by the authors.
  • Conflict of interest: The authors declare that they have no conflict of interest.
  • Availability of data and materials: No data was used in the preparation of this manuscript.

### v2

@@ 2.2. Zagreb indices of CCC-graphs of groups with a given central quotient
We have $\mathcal{CCC}(G)=K_{\frac{x(q-1)}{p}}\sqcup K_{x(p-1)}$, where $x=|Z(G)|$.Therefore, by Theorem 2.1, we get the expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$.
Note that $|v(\mathcal{CCC}(G))|=\frac{x(q-1)}{p}+x(p-1)=\frac{x}{p}(p^{2}+q-p-1)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{x(q-1)}{p}}{2}+\binom{x(p-1)}{2}=\frac{x}{2p^{2}}(p^{2}x+q^{2}x+p^{4}x+x-2qx-pq-2p^{3}x-p^{3}+p^{2}+p)$. It can be seen that
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{f(p,q,x)}{g(p,q,x)},$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+10p^{5}x^{2}-3p^{4}x^{3}-10p^{4}x^{2}+6p^{3}x^{2}-2p^{2}x^{2}$
and $g(p,q,x)=p^{2}(p^{2}+q-p-1)(p^{2}x+q^{2}x+p^{4}x+x-2qx-pq-2p^{3}x-p^{3}+p^{2}+p)$. On factorization, we have $f(p,q,x)=px^{2}(p-1)(q-1)(p+q-p^{2}-1)^{2}(p^{2}x+qx-x-2p-px)$. Suppose that $p^{2}x+qx-x-2p-px=p(x(p-1)-2)+x(q-1):=S_{1}+S_{2}$, where $S_{1}=p(x(p-1)-2)$ and $S_{2}=x(q-1)>0$ for all prime $q$ and $x\geq 1$. Clearly, $S_{1}\geq 0$ for all prime $p\geq 3$ and $x\geq 1$. Now, $f(2,q,1)=2(q-1)(q-3)^{2}(q-3)\geq 0$ for all $q\geq 3$. Also, $g(p,q,x)=\frac{2p^{5}}{x^{2}}\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. Therefore, $\frac{f(p,q,x)}{g(p,q,x)}\geq 0$ and hence the result follows. ∎
@@
In all the cases, we have $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}>\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ and hence the result follows. ∎
@@
Also, $|v(\mathcal{CCC}(G))|=\frac{x(p^{2}-1)}{p}+p^{2}\times\frac{x(p-1)}{p^{2}}=x(p-1)(p+2)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{x(p^{2}-1)}{p}}{2}+p^{2}\binom{\frac{x(p-1)}{p^{2}}}{2}=\frac{x(p-1)}{2p^{2}}(p^{3}x+p^{2}x-2p^{2}-2x-p)$. Therefore,
@@
$\left.\left.+p^{6}x^{2}(x-3)+3p^{5}x+p^{2}x(5x^{2}-4x-6)+p^{4}(4x^{2}+3x-2)+p^{3}(x^{3}+12x^{2}-1)\right)\right)$
@@
Also, $|v(\mathcal{CCC}(G))|=(p^{2}+p+1)\frac{x(p-1)}{p^{2}}=\frac{x(p^{3}-1)}{p^{2}}$ and $|e(\mathcal{CCC}(G))|=(p^{2}+p+1)\binom{\frac{x(p-1)}{p^{2}}}{2}$ $=\frac{x(p^{3}-1)}{2p^{4}}(px-x-p^{2})$. It can be seen that
$\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{p^{2}x^{2}-2px^{2}+x^{2}-2p^{3}x+2p^{2}x+p^{4}}{p^{4}}=\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}.$
@@
Also, $|v(\mathcal{CCC}(G))|=\frac{x(p^{2}-1)}{p}+kp\frac{x(p-1)}{p^{2}}+(p-k)\frac{x(p-1)}{p}=\frac{x(p-1)}{p^{2}}(2p^{2}+p)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{x(p^{2}-1)}{p}}{2}+kp\binom{\frac{x(p-1)}{p^{2}}}{2}+(p-k)\binom{\frac{x(p-1)}{p^{2}}}{2}=\frac{x(p-1)}{2p^{4}}(p^{5}x+2p^{4}x-2p^{3}x-2p^{4}-p^{2}x-p^{3}+2p^{2}kx-kpx-kp^{3}x)$. Therefore,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{f_{1}(p,k,x)}{g_{1}(p,k,x)},$
@@
Also, $|v(\mathcal{CCC}(G))|=(kp+1)\frac{x(p-1)}{p^{2}}+(p+1-k)\frac{x(p-1)}{p}=\frac{x(p-1)}{p^{2}}(p^{2}+p+1)$ and $|e(\mathcal{CCC}(G))|=(kp+1)\binom{\frac{x(p-1)}{p^{2}}}{2}+(p+1-k)\binom{\frac{x(p-1)}{p}}{2}=\frac{x(p-1)}{2p^{4}}(2kp^{2}x-kpx+px-x-p^{2}+p^{4}x+p^{4}-p^{2}x-p^{3}-kp^{3}x)$. Therefore,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{f_{2}(p,k,x)}{g_{2}(p,k,x)},$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+2kp^{3}x^{2}-2p^{3}x^{2}$
and $g_{2}(p,k,x)=p^{4}(p^{2}+p+1)(2kp^{2}x-kpx+px-x-p^{2}+p^{4}x+p^{4}-p^{2}x-p^{3}-kp^{3}x)$. On factorization, we have $f_{2}(p,k,x)=px^{2}(p-1)^{4}(p+1-k)(kp+1)(p^{2}x-x-2p^{2})$. For all $p\geq 2,1\leq k\leq p$ and $x\geq p^{2}$, we have $p+1-k>0$ and $x(p^{2}-1)>2p^{2}$. As such, $f_{2}(p,k,x)>0$. Also, $g_{2}(p,k,x)=\frac{2p^{10}}{x^{2}(p-1)^{2}}\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. Therefore, $\frac{f_{2}(p,k,x)}{g_{2}(p,k,x)}>0$.
@@ 3. Concluding remarks
Moreover, it was also seen that if $G$ is any non-abelian group of order $p^{3}$ or $p^{4}$, where $p$ is any prime, then $\mathcal{CCC}(G)$ satisfies Hansen-Vukičević conjecture. Using Theorem 2.8, Theorem 2.9 and the characterizations of $4$-centralizer (“$G$ is a $4$-centralizer finite group if and only if $\frac{G}{Z(G)}\cong\mathbb{Z}_{2}\times\mathbb{Z}_{2}$[3, Theorem 2]), $5$-centralizer (“$G$ is a $5$-centralizer finite group if and only if $\frac{G}{Z(G)}\cong\mathbb{Z}_{3}\times\mathbb{Z}_{3}$ or $D_{2\times 3}$[3, Theorem 4]) finite groups we have the following result.
@@
In addition, if $\Pr(G)=\frac{5}{14},\frac{2}{5},\frac{11}{27},\frac{1}{2}$ or $\frac{5}{8}$ then $\frac{G}{Z(G)}\cong D_{2\times 7},D_{2\times 5},D_{2\times 4},D_{2\times 3}$ or $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ (see [27, pp. 246] and [22, pp. 451]). Therefore, the following result follows in view of Theorem 2.8 and Theorem 2.9.
@@
It is worth noting that in this work we have not found any group $G$ such that $\mathcal{CCC}(G)$ does not satisfy Hansen-Vukičević conjecture. We conclude this paper with the following question.
@@
Acknowledgement. The first author gratefully acknowledges the Council of Scientific and Industrial Research for awarding the fellowship (File No. 09/0796(16521)/2023-EMR-I).
Conflict of interest. The authors declare that they have no conflict of interest.
Funding information. No funding was received by the authors.